OpenStudy (anonymous):

Solve for the variable L in this formula for math S=(n/2)(a+L)

OpenStudy (jkristia):

s(a+L) = (n/2) sa + sL = n/2 sL = (n/2) - sa then ....

OpenStudy (anonymous):

THEN???

OpenStudy (jkristia):

what is the next operation to isolate L on the left side?